OpenStudy (anonymous):

@amistre64 , @mathstudent55

OpenStudy (anonymous):

HELP!!!!!!!

OpenStudy (irishboy123):

what is cot(x)? it is a well known trig ratio - start by looking at something like this http://www.sosmath.com/trig/Trig5/trig5/trig5.html then you need to know what tan(x) is. it is opposite/adjacent. i can't vouch for the site but the formulae look good to me. in due course, you will know them all by heart:-)) tan = o/a sin = o/h cos = a/h some people here seem to remember the 3 basic trig ratios by the mnemonic "soh cah toa"

OpenStudy (anonymous):

it it would B???

OpenStudy (anonymous):

Or A

OpenStudy (anonymous):

@dan815 help!!!!

OpenStudy (dan815):

Tan theta = O/A u know that rightt

OpenStudy (dan815):